# Boundary Attention: Learning to Localize Boundaries under High Noise

### v2

@@ 4 Network Architecture
We seek a differential architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations. This bears similarities to the edge localization and edge-linking steps of early edge detectors like Canny’s, but with a more sophisticated local boundary model, and with mechanisms for spatial consistency that are learned instead of designed. It is also analogous to the original field of junctions algorithm [25], which uses coordinate descent for initialization and iterations of gradient descent for refinement. We want to replace both steps with something that is differentiable, faster, and able to scale to larger images.
@@
After considering a variety of alternatives, we settled on a particular family of architectures based on dot-product self-attention; and within this family, we performed an extensive set of ablations to determine the best performing model. We describe our final model here, and we provide the ablation details in Section 6. Importantly, all of our model’s components are invariant to discrete spatial shifts of the image, operating either on individual locations $k$, or on small neighborhoods of locations with spatially-shared parameters. This means that our model can be trained on small images and then deployed on much larger ones. Also, our model has only $207$k learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors. As a point of reference, the full EDTER model has 109M parameters [19].
@@
Our model is depicted in Figure 4. It implicitly uses higher-dimensional embeddings of the field elements, $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$. Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which, as shown in the supplement, provides enough capacity to learn a smooth latent representation of junction space. Use separate embeddings for the junction and windowing fields provides a disentangled representation of both.
@@
Given an input image, the network first applies a “neighborhood MLP-mixer”, which is a modified MLP-Mixer [24] that replaces the global spatial operations with convolutions of kernel size $3$. The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input. This block transforms the input image into a pixel-resolution map with ${D_{\gamma}}$ channels. We denote this by $\gamma_{0}[n]$ and refer to it as the initial “hidden state”. This hidden state is then refined by a sequence of eight boundary attention iterations, which we describe next. (See the supplement for a visualization of the decoded hidden states as they evolve.)
@@
The eight iterations of refinement are broken into two boundary attention blocks, each with learned weights. To process our input, we first add in a linear projection of the initial hidden state. This is essentially a skip connection that allows our network to retain information from the initial hidden state estimate at later stages of processing. Next, we copy our hidden state into two identical pieces. We concatenate a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other. We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy. We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even though global position cues are absent. We follow each self attention layer with a small MLP.
@@
To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping. We separate the windowing embedding (the last 8 dimensions) from the junction embedding (the first 64 dimensions) and project each through a linear layer. We map the hidden state to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$. These serve as the inputs to our gather and slice operators.
@@ 4.1 Training
We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end. Because junction space is non-convex and difficult to optimize over, this two step training procedure helps the network converge.
We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8. We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
We train our method using a combination of four global losses applied to global (i.e. sliced) fields, and two patch-wise losses applied to individual patches. The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
@@
where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as:
$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+(1-C),$ (11)
with $\beta$ and $C$ controlling how much weight to give pixels near boundaries. We set $\beta=0.1$, $\delta=1$, and anneal $C$ to give more weight to non-boundary locations later in training. We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement. Note that using noiseless feature maps for supervision in Equation 9 encourages the windowing functions to be large in smooth regions.
On top of the two supervision losses we apply two consistency losses from [25], that minimize the per-pixel variances $\nu_{f}[n]$ and $\nu_{d}[n]$. Similar to the supervision losses, we weigh those by $\alpha$ from Equation 11. These consistency losses encourage the junction shapes $\mathbf{g}$ in overlapping regions to agree. Minimizing $\nu_{f}[n]$ also encourages windowing functions to be large, because that increases the gather area, which in turn reduces noise in wedge features $\mathbf{f}_{nj}$ that are sliced to compute the variance $\nu_{f}[n]$ at each $n$.
@@
and we set $\delta^{\prime}=1$. These per-patch losses gives the network a more direct signal on how to adjust its weights than purely global losses, which average over multiple patches.
@@
Figure 6: Left: Qualitative comparison of our model versus other methods for noisy crops from the ELD dataset [26]. Our method provides more detail than other techniques at low noise and is more robust to high levels of noise. Right: Repeatability of the estimated boundaries of crops from the ELD dataset over increasing noise. Over multiple levels of real sensor noise, our method is the most consistent in predicting edges.
@@ 5 Experiments
We train our model on noisy synthetic data of colorful combinations of triangles and circles. We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training. To those crops we add Gaussian and/or Perlin noise [17], and with probability $0.1$ we average over the color channels to produce grayscale inputs. Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for validation. For a visualization of training data and training and optimization details, refer to the supplement.
@@
Since we focus on low-level edge detection, we primarily compare against other techniques with a similar focus and that also do not train on large semantic datasets: Canny [3], gPb [14], Structured Edges (SE) [7], and Field of Junctions (FoJ) [25]. However, since most recent work tends to focus on semantic edge detection, in the interest of being comprehensive, we also provide a comparison to several modern works: UAED [28], EDTER [19], and HED [27]. We note that this is not a clean comparison due to the difference in tasks – predicting edges with low-level cues vs predicting edges with semantic meaning.
Quantitative results and inference times across different noise levels.
For quantiative evaluation, our focus on low-level cues means that we cannot use semantic edge detection benchmarks like BSDS50 [2] since our method and other low-level baselines will both predict edges that are not semantically meaningful and fail to predict edges that are semantically meaningful but lack contrast. We instead rely on synthetic data where the ground truth edges can be determined with perfect precision and any level of noise can be added. In this experiment, we use geometric objects and per-pixel additive Gaussian noise with a random variance.
Figure S3 compares the performance and inference time of our method and baselines under different noise levels. The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches. Notably our method’s adaptive windowing function gives it at edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions. Our method is also orders of magnitude faster than FoJ as shown in the plot to the right.
@@
As shown in Figure 6, despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [26]. Our method produces crisp and well-defined boundaries despite high levels of noise. Figure 7 shows that our method makes reasonable boundary estimates for even complex in-the-wild images. Our supplement contains additional examples.
Figure 7: Our model can find compelling boundaries in natural in-the-wild images, despite training on simple synthetic data. Left image from the SIDD [1] dataset; right image taken with a phone camera.
@@
We quantify each model’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset. For each scene, we run the model on the lowest-noise image and then measure, via ODS F-score, how much its predictions change with increasing noise level. Figure 6 shows the averages of these scores across the collection. Our model provides more consistent results for increasingly noisy images than other methods.
6 Ablations
[…]
table:ablation1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2]. Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise. Our final model is boxed. It provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
[…]
@@ 6 Ablations
Loss Uniform $0.844$ $0.629$ $0.313$
Weighted* 0.872 0.673 0.348
Windowing Fixed - $3\times 3$ $0.834$ $0.546$ $0.205$
Fixed - $9\times 9$ 0.872 $0.662$ $0.328$
Fixed - $17\times 17$ $0.859$ $0.662$ $0.328$
@@
Table 2: Ablations: variations of our reported model. We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, the impact of a spatially weighted loss, and a fixed versus learned windowing function. Asterisks indicate our reported choices.
@@ 7 Conclusion
We have introduced a differentiable model that uses boundary attention to explicitly reason about geometric primitives such as edges, corners, junctions, and regions of uniform appearance in images. Our bottom-up, feedforward network can encode images of any resolution and aspect ratio into a field of geometric primitives that describe the local structure within every patch. Despite being trained on simple geometric synthetic data, our method can consistently localize boundaries under high levels of noise better than state-of-the-art deep networks, while achieving similar performance to recent optimization-based approaches in a fraction of the time.
Supplemental Material for Boundary Attention: Learning to Localize Boundaries under High Noise
@@
Another advantage of the present parameterization compared to that of the original [25] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}$ and $\mathbf{g}^{\prime}$. In our implementation we simply define interpolation variable $t\in[0,1]$ and compute interpolated junctions $\mathbf{g}^{(t)}=\{\boldsymbol{u}^{(t)},\theta^{(t)},\boldsymbol{\omega}^{(t)}\}$ using
$\boldsymbol{u}^{(t)}$ $=(1-t)\boldsymbol{u}+t\boldsymbol{u}^{\prime}$ (24)
$\tilde{\boldsymbol{\omega}}^{(t)}$ $=(1-t)\tilde{\boldsymbol{\omega}}+t\tilde{\boldsymbol{\omega}}^{\prime},$ (25)
@@
[…]
$(p,q)$ $=\operatorname{Slerp}\left((\cos\theta,\sin\theta),(\cos\theta^{\prime},\sin\theta^{\prime}),t\right),$
where $\operatorname{Slerp}()$ is the 2D geometric spherical linear interpolation operator,
$\operatorname{Slerp}(\boldsymbol{p},\boldsymbol{p}^{\prime},t)=\frac{\sin\left((1-t)\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}+\frac{\sin\left(t\Delta\theta\right)}{\sin\left(\Delta\theta\right)}\boldsymbol{p}^{\prime},$
with $\Delta\theta=\arccos(\boldsymbol{p}\cdot\boldsymbol{p}^{\prime})$. The bottom row of Figure 9 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
@@
Our dataset, which we call Kaleidoshapes, will be released upon publication, along with the code for generation, training and evaluation.
@@
Our model is designed to be purely local and bottom up, with all of its compositional elements operating on spatial neighborhoods in a manner that is invariant to discrete spatial shifts of an image. Its design also prioritizes having a small number of learnable parameters. Here we provide the details of the two blue blocks in the main paper’s Figure 3: Neighborhood MLP-Mixer and Neighborhood Cross-attention. We implement our model in JAX, and we will publicly share our model code and its pretrained weights upon publication.
@@
The neighborhood cross-attention block similarly enforces shift-invariance and weight sharing across spatial neighborhoods. Inside this block are two transformer layers whose cross-attention components are replaced with neighborhood cross-attention components that are restricted to a spatial neighborhood of pixels. We use $11\times 1$ 11 neighborhoods in our implementation. In each neighborhood containing a query token, we add a learned positional encoding to the key/value tokens which is relative to the neighborhood’s center and is the same for all neighborhoods. Then the query is updated using standard cross-attention with its neighborhood of key/values. We use 4 cross-attention heads. Like the standard transformer, each neighborhood cross attention component is followed by an MLP, dropout layer, and additive residual. To handle border effects, we zero-pad the key and value tokens so that every query attends to an $11\times 1$ 11 neighborhood, and then zero-out any attention weights involving zero-padded tokens.
@@
We pretrain the neighborhood MLP-mixer and the first boundary attention block on a simplified variation of our Kaleidoshapes dataset, where each image is $100\times 100$ and contains a single triangle and circle with additive zero-mean Gaussian noise. We omit the global losses of Equations $9$ and $10$ during this pretraining phase. This primes the network to learn meaningful hidden states $\boldsymbol{\gamma}[n]$ and prevents the “collapsing” of junctions, where the boundary-consistency loss (i.e. the sum over pixels of variance of distance $\nu_{d}[n]$) dominates and the network learns to predict all-boundaryless patches that are globally consistent but inaccurate. Because of data imbalance—only a small fraction of regions $\Omega_{n}(x)$ contain corners or junctions—we add an additional spatial importance mask to prioritize the regions that contain a corner (i.e., a visible triangle vertex) or a junction (i.e., an intersection between a circle and a triangle’s edge). Our data generation process produces a list of all non-occluded vertices and intersections in each image, and we use these values to create a spatial importance mask with gaussians centered at each of these points. In practice, we use gaussians with a standard deviation of 7 pixels. This mask is added to the loss constant $C$.
The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block. We use $100k$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block. Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes). Our trained weights will be released upon publication.
@@
S4.1  Generalization to Unseen Shapes
During training, our model observes only circles and triangles of varying sizes and colors. We measure how well our model generalizes to unseen shapes by creating a new synthetic dataset containing rectangles, polygons, and ellipses of various sizes, colors, and rotations. We evaluate our model against low-level methods as well as methods trained on semantic data, and note the results here closely align with the performance trends on kaleidoshapes (Fig. 5), suggesting that our bottom-up model can generalize to unseen geometries.
Fig. S3: We measure how well our model can generalize to shapes outside of the training distribution by evaluating our model and others on noisy synthetic data composed exclusively of unseen shapes.
S4.2  Linear Interpolation in Junction Space
Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state. Figure S4 visualizes some properties of its learned embedding of junction space. We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$. We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center. For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
Fig. S4: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$. Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$. The embedding has learned to be smooth and have an intuitive zero.
S4.3  Evolution of the Outputs
Figure S5 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement. Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, which we then apply a non-linearity to in order to amplify the less prominent boundaries. We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
Fig. S5: Evolution of boundaries during iterations, in reading order. Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
S4.4  Qualitative Results for Natural Images

### v3

@@ 4 Network Architecture
We seek a differentiable architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations. This approach is motivated by the edge localization and edge-linking steps of early edge detectors like Canny’s, but has a more sophisticated local boundary model, and mechanisms for spatial consistency that are learned instead of designed. It is also analogous to the original field of junctions algorithm [23], which uses coordinate descent for initialization and iterations of gradient descent for refinement. We want to replace both steps with something that is differentiable, faster, and able to scale to larger images.
@@
After considering a variety of alternatives, we settled on a particular family of architectures based on dot-product self-attention; and within this family, we performed an extensive set of ablations to determine the best performing model. We describe our final model here, and we provide the ablation details in Section 6. Importantly, all of our model’s components are invariant to discrete spatial shifts of the image, operating either on individual locations $k$, or on small neighborhoods of locations with spatially-shared parameters. This means that our model can be trained on small images and then deployed on much larger ones. Also, our model has only $207,000$ learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors. As a point of reference, Diffusion Edge, a recent diffusion-based semantic boundary detection model uses on the order of 300 million parameters [26].
@@
Our model is depicted in Figure 4. It represents the field elements as higher-dimensional embeddings $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$. Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which we show in our experiments provides enough capacity to learn a smooth latent representation of junction space. Using separate embeddings for the junction and windowing fields provides a disentangled representation of both.
@@
Given an input image, the network first applies a “neighborhood MLP-mixer”, which is a modified MLP-Mixer [22] that replaces the global spatial operations with convolutions of kernel size $3$. The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input. This block transforms the input image into a pixel-resolution map with ${D_{\gamma}}$ channels. We denote this by $\gamma_{0}[n]$ and refer to it as the initial “hidden state”. This hidden state is then refined by a sequence of eight boundary attention iterations, which we describe next. (See our experiments for a visualization of the decoded hidden states as they evolve.)
@@
The eight iterations of refinement are broken into two blocks, each with learned weights. In each iteration, we first add a linear mapping of the initial hidden state to the current hidden state, which acts as a skip connection. Next, we clone our hidden state, concatenating a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other. We treat the copies as the inputs to neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy. We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even though global position cues are absent. We follow each self attention layer with a small MLP.
@@
To transform our output or intermediary hidden state into junction space and render patch visualizations, we use a simple linear mapping. We separate the windowing embedding (the last 8 dimensions) from the junction embedding (the first 64 dimensions) and project each through a linear layer. We map the junction embedding to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$. These serve as the inputs to our gather and slice operators.
@@ 4.1 Training
Estimating the best junctions for noisy image patches is a non-convex optimization that is prone to getting stuck in local minima [23]. We find that training our network in three stages with increasingly complex synthetic training data produces the best model. First we train the network on small $21\times 21$ images containing a single junction corrupted by low amounts of Gaussian noise. Upon convergence, we retrain the network on larger $100\times 100$ images containing a single circle and triangle corrupted by moderate Gaussian noise. Finally, we retrain the network on a high-noise synthetic dataset containing $125\times 125$ images consisting of many overlapping triangles and circles. (See supplement for examples.)
We calculate our loss using the outputs of the two final iterations of our network, where the loss of the final output is weighted three times as much as the loss applied to the output before it. This encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
We train our model using a combination of four global losses applied to global (i.e. sliced) fields, and two patch-wise losses applied to individual patches. The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and distance maps:
@@
where $\mathbf{f}_{\text{GT}}$ and $d_{\text{GT}}$ are the ground truth features and distance maps, respectively, and $\alpha[n]$ is a pixel importance function defined as
$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+C,$ (11)
with $\beta$ and $C$ controlling how much weight to give pixels near boundaries. We set $\beta=0.1$, $\delta=1$, and $C=0.3$. (The supplement contains additional tests with a more involved importance mask.) Using noiseless feature maps for supervision in Equation 9 has the effect of encouraging windowing functions to be as large as possible, because larger regions imply averaging over a greater number of noisy input pixel values.
On top of the two supervision losses we apply two consistency losses borrowed from [23], which minimize the per-pixel variances $\nu_{f}[n]$ and $\nu_{b}[n]$. Here, we weight them by $\alpha[n]$ from Equation 11. These consistency losses encourage the junction shapes $\mathbf{g}$ in overlapping regions to agree. Minimizing $\nu_{f}[n]$ also provides a second mechanism to encourage windowing functions to be large. Larger windows increase gather area, thereby reducing noise in wedge features $\mathbf{f}_{nj}$ that are sliced to compute variance $\nu_{f}[n]$ at each pixel $n$.
@@
with $\delta^{\prime}=1$. These per-patch losses provide a more direct signal for adjusting model weights, compared to per-pixel losses which average over multiple patches.
@@
Figure 6: Left: Qualitative comparison of our model versus other methods for a noisy crop from the ELD dataset [24]. See the supplement for more examples. Our method provides more detail than other techniques at low noise and is more robust to high levels of noise. Right: Repeatability of the estimated boundaries of crops from the ELD dataset over increasing noise. Over multiple levels of real sensor noise, our method is the most consistent in predicting edges.
@@ 5 Experiments
In the final stage of training, we use noisy synthetic data of many randomly-colored overlapping triangles and circles. We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training. To those crops we add Gaussian or Perlin noise [16], and with probability $0.1$ we average over the color channels to produce grayscale inputs. Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for validation.
@@
Since we focus on bottom-up edge and corner detection, we compare against other techniques with a similar focus and that, like us, do not train on large semantic datasets: Canny [2], gPb [13], Structured Edges (SE) [6], and Field of Junctions (FoJ) [23].
Quantitative results.
For quantitative evaluation we cannot use semantic edge detection benchmarks like BSDS500 [1], because our model and the comparisons are designed to predict all edges, including those that are not semantically meaningful. We instead rely on synthetic data, where the ground truth edges can be determined with perfect precision, and inputs can be controllably noised. Our evaluation data comprises geometric objects and per-pixel additive Gaussian noise with a random variance.
Figure 5 compares the performance and inference time of our method and baselines under different noise levels. The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches. Notably our method’s adaptive windowing function gives it an edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions. Our method is also orders of magnitude faster than FoJ, as shown at right.
@@
As shown in Figure 6, despite being trained on synthetic data, our method can detect edges in real photographs with multiple levels of real sensor noise present in ELD [24]. Our method produces crisp and well-defined boundaries despite high levels of noise. The supplement includes additional examples that show that our method makes reasonable boundary estimates for other captured images.
@@
We quantify each method’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset. For each scene, we run the model on the lowest-noise image and then measure, via ODS F-score, how much its predictions change with increasing noise level. Figure 6 shows the averages of these scores across the collection. Our model provides more consistent results for increasingly noisy images than other methods, in addition to capturing fine details that other methods miss.
Properties of learned embedding.
We find that our model learns a spatially smooth embedding $\boldsymbol{\gamma}\in\mathbb{R}^{D_{\gamma}}$ of junction space $\mathbf{g}\in{\cal G}$. In Figure 7 we generate equally-spaced samples $\boldsymbol{\gamma}_{i}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to $\boldsymbol{0}$ and then to a particular $\boldsymbol{\gamma}_{b}$; and then to each sample we apply the learned embedding to compute and visualize the implied junction $\mathbf{g}_{i}$. We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center. For visual comparison, we show an analogous geometric interpolation in junction space ${\cal G}$ (see the supplement for expressions) from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$.
Figure 7: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$. Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$. The embedding has learned to be smooth and have an intuitive zero.
Figure 8: Evolution of boundaries during iterations, in reading order. Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
@@
Evolution.
Figure 8 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement. Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, to which we apply a non-linearity that amplifies less-prominent boundaries. We see that early iterations are exploratory and unstructured, and that later iterations agree.
6 Ablations
Table 1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2]. Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise. Our final model (boxed) provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
Table 2 compares many other variations of the boxed model from Tab. 1, with asterisks indicating the model specifications used in the paper. Iterations varies the number of attention-iterations within each block, and Neighborhood varies the attention neighborhood size. MLP Input varies the features concatenated to the hidden state prior to the pre-attention MLP and shows that replacing the gathered colors $\bar{\mathbf{f}}^{t}$ with a constant array or the input image values $\mathbf{f}$ performs worse. Windowing shows that using fixed, square windowing functions $w_{n}(x)$ of any size performs equal (at low noise) or worse than inferring them adaptively. Somewhat intuitively, a small $9\times 9$ patch size performs well at low noise, but performance lags under noisy conditions where using a larger patch size increases the spatial extent of the communication across patches.
@@ 6 Ablations
Windowing Fixed 0.872 $0.662$ $0.328$
@@
Table 2: Ablations: variations of our reported model. We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, and a fixed versus learned windowing function. Asterisks indicate our reported choices.
@@ 7 Conclusion
We have introduced a differentiable model that uses boundary attention to explicitly reason about geometric primitives such as edges, corners, junctions, and regions of uniform appearance in images. Our bottom-up, feedforward network can encode images of any resolution and aspect ratio into a field of geometric primitives that describe the local structure within every patch. This work represents a step to the goal of synergizing the benefits of low-level parametric modeling with the efficiency and representational power of deep learning.
Supplemental Material for Boundary Attention: Learning curves, corners, junctions and grouping
@@
Another advantage of the present parameterization compared to that of the original [23] is that it is a simply-connected topological space and so allows for defining mechanisms for smoothly interpolating between any two junctions $\mathbf{g}=\left(\boldsymbol{u},\theta,\boldsymbol{\omega}\right)$ and $\mathbf{g}^{\prime}=\left(\boldsymbol{u}^{\prime},\theta^{\prime},\boldsymbol{\omega}^{\prime}\right)$. In our implementation we define interpolation variable $t\in[0,1]$ and compute interpolated junctions $\mathbf{g}(t)=\{\boldsymbol{u}(t),\theta(t),\boldsymbol{\omega}(t)\}$ using a simple combination of linear and spherical linear interpolation:
$\boldsymbol{u}(t)$ $=(1-t)\boldsymbol{u}+t\boldsymbol{u}^{\prime}$ (24)
$\tilde{\boldsymbol{\omega}}(t)$ $=(1-t)\tilde{\boldsymbol{\omega}}+t\tilde{\boldsymbol{\omega}}^{\prime},$ (25)
@@
$\theta(t)=\theta+t\Delta\theta,$ (26)
with
$\Delta\theta=\begin{cases}\theta^{\prime}-\theta-2\pi,&\text{if }\theta^{\prime}-\theta>\pi\\ \theta^{\prime}-\theta+2\pi,&\text{if }\theta^{\prime}-\theta<-\pi\\ \theta^{\prime}-\theta,&\text{otherwise},\end{cases}$
assuming $\theta,\theta^{\prime}\in[0,2\pi)$. The bottom row of Figure 7 in the main paper visualizes a set of samples from smooth trajectories in junction space using this mechanism.
@@
Our dataset, which we call Kaleidoshapes, is available publicly, along with the code for generation, training and evaluation.
@@
Our model is designed to be purely local and bottom up, with all of its compositional elements operating on spatial neighborhoods in a manner that is invariant to discrete spatial shifts of an image. Its design also prioritizes having a small number of learnable parameters. Here we provide the details of the two blue blocks in the main paper’s Figure 3: Neighborhood MLP-Mixer and Neighborhood Cross-attention. Our model was implemented with JAX and its code and trained weights are available publicly.
@@
The neighborhood cross-attention block similarly enforces shift-invariance and weight sharing across spatial neighborhoods. Inside this block are two transformer layers whose cross-attention components are replaced with neighborhood cross-attention components that are restricted to a spatial neighborhood of pixels. We use $11\times 1$ 11 neighborhoods in our implementation, which our ablations showed produced the best performance. In each neighborhood containing a query token, we add a learned positional encoding to the key/value tokens which is relative to the neighborhood’s center and is the same for all neighborhoods. Then the query is updated using standard cross-attention with its neighborhood of key/values. We use 4 cross-attention heads. Like the standard transformer, each neighborhood cross attention component is followed by an MLP, dropout layer, and additive residual. To handle border effects, we zero-pad the key and value tokens so that every query attends to an $11\times 1$ 11 neighborhood, and then zero-out any attention weights involving zero-padded tokens.
@@
During the first two stages of training where we train our model on junction images and simplified circle triangle images, we omit the global losses of Equations $9$ and $10$. This primes the network to learn meaningful hidden states $\boldsymbol{\gamma}[n]$ and prevents the “collapsing” of junctions, where the boundary-consistency loss (i.e. the sum over pixels of variance of distance $\nu_{d}[n]$) dominates and the network learns to predict all-boundaryless patches that are globally consistent but inaccurate. Because of data imbalance—only a small fraction of regions $\Omega_{n}(x)$ contain corners or junctions—we add an additional spatial importance mask to prioritize the regions that contain a corner (i.e., a visible triangle vertex) or a junction (i.e., an intersection between a circle and a triangle’s edge). Our data generation process produces a list of all non-occluded vertices and intersections in each image, and we use these values to create a spatial importance mask with gaussians centered at each of these points. In practice, we use gaussians with a standard deviation of 7 pixels. This mask is added to the loss constant $C$.
The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block. We use $100k$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block. Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes).
@@
S4.1  Qualitative Results for Natural Images